OpenStudy (anonymous):

Three party-goers are in the corner of the ballroom having an intense argument. You walk over to settle the debate. They are discussing a function g(x). You take out your notepad and jot down their statements. Professor McCoy: She says that 2 is a zero of g(x) because long division with (x + 2) results in a remainder of 0. Ms. Guerra: She says that 2 is a zero of g(x) because g(2) = 0. Mr. Romano: He says that 2 is a zero of g(x) because synthetic division with 2 results in a remainder of 0.

OpenStudy (anonymous):

Correct the reasoning of any inaccurate reasoning by the party-goers in full and complete sentences. Make sure you reference any theorems that support your justifications.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

let's make up a simple polynomial that has a zero at x=2 for example: g(x)= x-2 when does g(x)=0 ? when x- 2 = 0 or x= 2 now test the first statement long division with (x + 2) results in a remainder of 0. (x-2)/(x+2) what do you get ?

OpenStudy (anonymous):

x^2+2x-2x-4 x^2-4

OpenStudy (phi):

when you divide, you get something smaller. But the important point here is it does not go evenly. statement one is wrong.

OpenStudy (phi):

statement 2 is the definition of a zero, so it is correct. to test the last statement, try synthetic division 2 | 1 -2 |____2__ 1 0

OpenStudy (anonymous):

Oh, I did FOIL. I'm not sure how to divide them

OpenStudy (phi):

when you have time, here is how to do it http://www.khanacademy.org/math/algebra/multiplying-factoring-expression/dividing_polynomials/v/polynomial-division but it does not divide evenly btw, in the same play list is a video on synthetic division.

OpenStudy (anonymous):

I've done synthetic dvision

OpenStudy (anonymous):

What about Ms. Guerra? She says that 2 is a zero of g(x) because g(2) = 0.

OpenStudy (anonymous):

@phi ?

OpenStudy (phi):

I told you up above. statement 2 is the definition of a zero, so it is correct.

OpenStudy (anonymous):

Oh I'm sorry. I missed that post. lol

OpenStudy (anonymous):

So is the last statement correct since it equals zero? 2 | 1 -2 |____2__ 1 0

OpenStudy (anonymous):

@phi ...

OpenStudy (anonymous):

Are their steps I need to take after that?